# Atomic Josephson Parametric Amplifier

### v2

@@ Abstract
We study the dynamics of a driven atomic Josephson junction that we propose as a parametric amplifier. By periodically modulating the position of the barrier, we induce a small current across the junction, serving as our input signal. The pump field is implemented by modulating the barrier height at twice the Josephson plasma frequency. The resulting dynamics exhibit parametric amplification of the signal through nonlinear mixing between the signal and pump fields, which is reflected by specific patterns of density waves and phase excitations. We employ a classical-field dynamics method and benchmark our results against driven circuit dynamics. This work paves the way for tunable quantum amplifiers in atomtronic circuits, with potential applications in several fields including precision measurements and quantum information processing.
@@
Figure 1: Atomic Josephson parametric amplifier (JPA). (a) Simulation of a Josephson junction, which consists of two 2D clouds separated by a Gaussian tunnel barrier of height $V_{0}$ and width $w$ (indicated by two dotted vertical lines). We use $V_{0}/\mu=1.5$ and $w=1\,\mu\mathrm{m}$, where $\mu$ is the mean-field energy. (b) Sketch of the JPA protocol. We periodically modulate the barrier position using $x(t)=x_{0}+x_{1}\sin(2\pi f_{\mathrm{s}}t)$, where $x_{1}$ is the amplitude and $f_{\mathrm{s}}$ is the frequency. For the pump we modulate the barrier height, i.e., $V(t)=V_{0}\bigl{(}1+A_{\mathrm{d}}\cos(2\pi f_{\mathrm{d}}t)\bigr{)}$, where $A_{\mathrm{d}}$ is the amplitude and $f_{\mathrm{d}}$ is the frequency. (c) Sketch of the operation of a JPA. We modulate the barrier height near twice the plasma frequency $f_{0}$ of the JJ. The signal mode at $f_{\mathrm{d}}/2+\delta f$ is amplified and an idler mode at frequency $f_{\mathrm{d}}/2-\delta f$ is created, with $\delta f$ being the tuning parameter.
@@
Figure 3: Nonlinear mixing in the power spectrum $S_{\omega}=|z(\omega)|^{2}$ for signal frequencies between $f_{\mathrm{s}}=28$ and $52\,\mathrm{Hz}$ at $\tilde{I}_{1}=0.42$. For each column the spectrum is normalized by its maximum value $S_{\omega,\mathrm{m}}$. (a) Without pump the spectrum mainly displays the central peak at $f_{\mathrm{s}}=f$. (b , c) In the presence of pump there is another peak at $f_{\mathrm{d}}-f_{\mathrm{s}}$ (indicated by white dashed line), located at the frequency of the signal mirrored at the plasma frequency $f_{0}$. The case of $f_{\mathrm{s}}=f_{0}$, where signal and idler are at the same frequency, is a special case known as the degenerate case. (d) The result of the RSJ circuit model is shown for the same system parameters and $A_{\mathrm{d}}=0.5$; see text.
@@
which include the barrier potential given by $V({\bf r},t)=V_{0}(t)\exp\bigl{[}-2\bigl{(}x-x(t)\bigr{)}^{2}/w^{2}\bigr{]}$. $V_{0}(t)$, $w$ and $x(t)$ are the barrier’s strength, width and location. We perform numerical simulations by discretizing space on a lattice of size $N_{x}\times N_{y}=100\times 200$ with the discretization length $l=0.5\,\mu\mathrm{m}$. To illustrate the key idea, we focus on a concrete realization using ${}^{6}\mathrm{Li}_{2}$ molecules, though our protocol is applicable to any cold-atom degenerate gas. We set the density to $n\approx 5.6\,\mu\mathrm{m}^{-2}$, interaction parameter $\tilde{g}=0.1$, and temperature ratio $T/T_{0}=0.06$, with system dimensions $L_{x}\times L_{y}=50\times 100\,\mu\mathrm{m}^{2}$. The critical temperature $T_{0}$ is estimated by $T_{0}=2\pi n\hbar^{2}/(mk_{\mathrm{B}}{\cal{D}}_{c})$, where ${\cal{D}}_{c}=\ln(380/\tilde{g})$ is the critical phase-space density [63, 64]. We use $w/\xi=1.1$ and $V_{0}/\mu\equiv\tilde{V}_{0}=1.5$, where $\xi=\hbar/\sqrt{2mgn}$ is the healing length and $\mu=gn$ is the mean-field energy. To create the weak link at $x(t)=x_{0}=L_{x}/2$, we linearly ramp up $V_{0}$ over $200\,\mathrm{ms}$ and then wait for $50\,\mathrm{ms}$. The density distribution $n(x,y)=|\psi({\bf r},t)|^{2}$ is shown in Fig. 1(a), which is averaged over the initial ensemble. For the signal we periodically modulate the barrier position using $x(t)=x_{0}+x_{1}\sin(2\pi f_{\mathrm{s}}t)$, where $x_{1}$ is the amplitude and $f_{\mathrm{s}}$ is the frequency, see Fig. 1(b). The resulting barrier velocity $v(t)=v_{1}\cos(2\pi f_{\mathrm{s}}t)$, with $v_{1}=2\pi f_{\mathrm{s}}x_{1}$, induces an oscillating current across the junction. We calculate the atom number $N_{L}(t)$ ($N_{R}(t)$) in the left (right) reservoir to determine the imbalance $z(t)=(N_{L}(t)-N_{R}(t))/N$, where $N=(N_{L}+N_{R})$ is the total atom number. For the spectrum analysis we calculate the power spectrum
@@
[…]
Characterization of the JPA dynamics.— To characterize the dynamical regimes we map out the imbalance spectrum for signal frequencies between $f_{\mathrm{s}}=20$ and $100\,\mathrm{Hz}$, which exhibits JPA resonances and other excitation peaks [65]. Fig. 3 highlights the interplay between the signal and idler modes for the lowest JPA resonance. The mode weights are controlled by the pump amplitude. To describe the JPA dynamics we introduce a driven RSJ (resistively shunted junction) circuit model. The barrier velocity results in a modulation of the current across the junction given by $I_{1}\cos(\omega_{s}t)$, where $I_{1}=v_{1}I_{c}/v_{c}$ in which $v_{c}$ is the critical velocity associated with the critical current $I_{c}$ [65]. The barrier height modulation changes the tunneling energy, which can be incorporated as $I_{c}\bigl{(}1+i_{\mathrm{d}}\sin(\omega_{\mathrm{d}}t)\bigr{)}$. The effective pump amplitude $i_{\mathrm{d}}$ is linked to $A_{\mathrm{d}}$ based on the prediction of the critical current $I_{c,\mathrm{p}}=I_{\text{b}}t_{0}(\tilde{V}_{0},w)$ derived for an ideal Josephson junction [66]. The bulk current $I_{\text{b}}=cn_{0}L_{y}$ is determined using the sound velocity $c=\sqrt{gn/m}$ and the condensate density $n_{0}$. We use the variational solution of the tunneling amplitude $t_{0}(\tilde{V},d)$ obtained across a rectangular barrier of width $d$ and height $\tilde{V}$ [66], with $d\approx 1.1w$ and $\tilde{V}=\tilde{V}_{0}$ [65]. The Kirchhoff law of the driven RSJ circuit reads
$I_{1}\cos(\omega_{s}t)=I_{c}\bigl{(}1+i_{\mathrm{d}}\sin(\omega_{\mathrm{d}}t)\bigr{)}\sin\phi-G\Delta\mu,$ (4)
where $\phi=\phi_{L}-\phi_{R}$ is the phase difference across the junction and $G$ is the conductance. The Josephson relation for the phase dynamics is $\hbar\dot{\phi}=-\Delta\mu$, with $\Delta\mu$ playing the role of the voltage across the junction. We numerically solve Eq. 4 for various signal and pump parameters and obtain the time evolution $\Delta\mu(t)$. The values of $I_{c}=135\,(10^{3}/\mathrm{s})$ and $\hbar G=59.8$ are determined from the simulated current-chemical potential relation [65]. With $I_{c}$ and the effective charging energy $E_{C}=4(\partial\mu/\partial N)$ we estimate the Josephson frequency $\omega_{\text{J}}=\sqrt{I_{c}E_{C}/\hbar}$ and obtain $f_{\text{J}}=38\,\mathrm{Hz}$ in agreement with the value of $f_{0}$. The result of the spectrum $|\Delta\mu(\omega)|^{2}$ is shown in Fig. 3d, which captures the dynamical response of the JPA. We note that the circuit model also describes higher JPA resonance and other excitation peaks well [65].
To quantify the overall amplification, we define a dimensionless gain $G_{\mathrm{s}}=S_{\omega}(A_{\mathrm{d}})/S_{\omega}(0)-1$, where $S_{\omega}(A_{\mathrm{d}})$ represents the signal peak value and $S_{\omega}(0)$ corresponds to the result without pump. In Fig. 4(a) we demonstrate that $G_{\mathrm{s}}$ initially increases approximately quadratically with $A_{\mathrm{d}}$, and then reaches a maximum for a value of $A_{\mathrm{d}}$, that decreases with increasing $\tilde{I}_{1}=I_{1}/I_{c}$. The gain $G_{\mathrm{s}}$ is exemplified in Fig. 4(c) in frequency space. For larger values of $A_{\mathrm{d}}$, the gain $G_{\mathrm{s}}$ decreases. While the nonlinear increase is well captured by the circuit model, the model fails to describe the decrease beyond the maximum, primarily visible for larger $\tilde{I}_{1}$. Fig. 4(b) presents the signal-to-noise ratio (SNR), defined as $\bar{S}_{\omega}/\Delta\bar{S}_{\omega}$, where we compute the mean signal value $\bar{S}_{\omega}=\left\langle S_{\omega}(A_{\mathrm{d}})-S_{\omega}(0)\right\rangle$ and the noise $\Delta\bar{S}_{\omega}$ from the variance determined by $\Delta\bar{S}_{\omega}^{2}=\left\langle(S_{\omega}(A_{\mathrm{d}})-S_{\omega}(0)-\bar{S}_{\omega})^{2}\right\rangle$. Here, .$\langle$.$\rangle$ . . ⟩ denotes a statistical average over $100$ samples of the ensemble. The SNR depends on both $A_{\mathrm{d}}$ and $\tilde{I}_{1}$, with a visible increase with increasing $A_{\mathrm{d}}$. The maximal regime of the SNR is in the same regime as the maximal gain $G_{\mathrm{s}}$, suggesting that this is the optimal regime of operation.
@@
Figure 4: JPA characteristics. (a) Dimensionless gain $G_{\mathrm{s}}=S_{\omega}(A_{\mathrm{d}})/S_{\omega}(0)-1$ as a function of $A_{\mathrm{d}}$ is shown for signal amplitudes between $\tilde{I}_{1}=0.12$ and $0.42$. The results of the RSJ model are indicated by the dashed lines. (b) Signal-to-noise ratio (SNR) corresponding to the results shown in (a). (c) Signal peak in the power spectrum $S_{\omega}=|z(\omega)|^{2}$ at $\tilde{I}_{1}=0.42$ is shown for different pump amplitudes $A_{\mathrm{d}}$.
To understand the underlying mechanism of the JPA we examine the time evolution of the density and the phase at the junction in Fig. 5. Without pump, the barrier velocity modulation results in a smooth excitation of the density, which is created by the change of junction phase that is synchronized with the signal frequency. When the pump field is on, there are higher-order density excitations that result from the modulation of the phase by the pump field. These can be explained with a perturbative solution of the equation $\tilde{I}_{1}\cos(\omega_{\mathrm{s}}t)=\bigl{(}1+A_{\mathrm{d}}\sin(\omega_{\mathrm{d}}t)\bigr{)}\sin\phi$, assuming a small modulation parameter $A_{\mathrm{d}}$ and expanding $\phi$ as $\phi=\phi_{0}+A_{\mathrm{d}}\phi_{1}+\mathcal{O}(A_{\mathrm{d}}^{2})$. At zeroth order we obtain $\phi_{0}\approx\tilde{I}_{1}\cos(\omega_{\mathrm{s}}t)$, which explains the smooth oscillation of the phase without the pump. The first-order term gives $\phi_{1}\approx-\tilde{I}_{1}\cos(\omega_{\mathrm{s}}t)\sin(\omega_{\mathrm{d}}t)\bigr{)}$, highlighting the intrinsic nonlinear mixing between the pump and signal field. In Fig. 5(d), the perturbative solutions consistently capture the dynamics of our atomic JPA.
@@
Figure 5: Underlying mechanism. (a) Time evolution of the density $\delta n(x,t)=n({\bf r},t)-n({\bf r},0)$, averaged over the $y$ direction and the ensemble, is shown for $\tilde{I}_{1}=0.21$ without the pump. (b) Dynamics in the presence of pump with $A_{\mathrm{d}}=0.8$. (c) Ensemble averaged phase difference $\phi$ at the junction for the cases in (a) and (b). (d) The results of zeroth and first order expansion of Eq. 4 without the resistive term; see text.
@@
Conclusions and outlook.— We have studied the dynamics of a driven atomic Josephson junction (JJ) using classical-field simulations. The JJ is formed by coupling two two-dimensional atom clouds with a tunnel barrier, where we periodically modulate both the position (input signal) and the height (pump) of the barrier. By tuning the pump frequency to twice the Josephson plasma frequency, we have demonstrated parametric amplification for the signal. Our system works in the $\mathrm{Hz}$ frequency range, that is beyond the operating range of the superconducting JPAs. The current cold-atoms technology allows to accurately image the microscopic dynamics of the system [57, 58]. In this case, we have an unprecedented access to the density and phase dynamics at the junction responsible for the amplification and the intrinsic nonlinear mixing between the signal and pump fields. Our results can be experimentally probed using in-situ imaging and matter-wave interferometry, where the former enables measurement of the density imbalance and the latter allows detection of the underlying JPA mechanism. Our proposed protocol can be implemented in driven Josephson junctions experimentally realized in [57, 58].
@@
Our findings have broad implications for the understanding of the basic mechanism on which parametric amplification relies in quantum systems. For the specific case of superconducting JPAs, the microscopic phase-density dynamics in the system, can be probed by suitable Josephson current measurements.

### v3

@@ Abstract
We study the dynamics of a driven atomic Josephson junction that we propose as a parametric amplifier. By periodically modulating the position of the barrier, we induce a small current across the junction, serving as our input signal. The pump field is implemented by modulating the barrier height at twice the Josephson plasma frequency. The resulting dynamics exhibit parametric amplification of the signal through nonlinear mixing between the signal and pump fields, which is encoded in a specific microscopic pattern of density waves and phase excitations that can be addressed within the experimental cold atoms capabilities. This work paves the way for tunable amplifiers in atomtronic circuits, with potential applications in several fields including precision measurements and quantum information processing. At the same time, our analysis provides the microscopic explanation of the general notion of parametric amplification occurring in nonlinear coherent devices.
@@
Figure 1: Atomic Josephson parametric amplifier (JPA). (a) Simulation of a Josephson junction, which consists of two 2D clouds separated by a Gaussian tunnel barrier of height $V_{0}$ and width $w$ (indicated by two dotted vertical lines). We use $V_{0}/\mu=1.5$ and $w/\xi=1.1$, where $\mu$ is the mean-field energy and $\xi$ is the healing length. (b) Sketch of the JPA protocol. We periodically modulate the barrier position using $x(t)=x_{0}+x_{1}\sin(2\pi f_{\mathrm{s}}t)$, where $x_{1}$ is the amplitude and $f_{\mathrm{s}}$ is the frequency. For the pump we modulate the barrier height, i.e., $V(t)=V_{0}\bigl(1+A_{\mathrm{d}}\cos(2\pi f_{\mathrm{d}}t)\bigr)$, where $A_{\mathrm{d}}$ is the amplitude and $f_{\mathrm{d}}$ is the frequency. (c) Sketch of the operation of a JPA. We modulate the barrier height near twice the plasma frequency $f_{0}$ of the $J$. The signal mode at $f_{\mathrm{d}}/2+\delta f$ is amplified and an idler mode at frequency $f_{\mathrm{d}}/2-\delta f$ is created, with $\delta f$ being the tuning parameter.
@@
Figure 3: Nonlinear mixing in the power spectrum $S_{\omega}=|z(\omega)|^{2}$ for signal frequencies between $f_{\mathrm{s}}=28$ and $52\,\mathrm{Hz}$ at $\tilde{I}_{1}\equiv I_{1}/I_{c}=0.42$. For each column the spectrum is normalized by its maximum value $S_{\omega,\mathrm{m}}$. (a) Without pump the spectrum mainly displays the central peak at $f_{\mathrm{s}}=f$. (b, c) In the presence of pump there is another peak at $f_{\mathrm{d}}-f_{\mathrm{s}}$ (indicated by white dashed line), located at the frequency of the signal mirrored at the plasma frequency $f_{0}$. The case of $f_{\mathrm{s}}=f_{0}$, where signal and idler are at the same frequency, is a special case known as the degenerate case. (d) The result of the RSJ circuit model for $\tilde{I}_{1}=0.42$ and $A_{\mathrm{d}}=0.8$; see text.
@@
[…]
@@
where $z(\omega)=(1/\sqrt{T_{n}})\int dt\,z(t)\exp(-i\omega t)$ is the Fourier transform of $z(t)$. $T_{n}$ is the sampling time for the numerical Fourier transform. The junction displays a well-defined resonance at the Josephson plasma frequency $f_{0}$, which we determine by analyzing $S_{\omega}$ at varying frequency, obtaining $f_{0}=39.4\,\mathrm{Hz}$ in the limit of $v_{1}\rightarrow 0$ (see End Matter for details). For the parametric pump we modulate the barrier height, i.e., $V(t)=V_{0}\bigl(1+A_{\mathrm{d}}\cos(2\pi f_{\mathrm{d}}t)\bigr)$, where $A_{\mathrm{d}}$ is the amplitude and $f_{\mathrm{d}}$ is the frequency, as depicted in Fig. 1(b). This alters the nonlinear inductance of the junction, which gives rise to amplification of the signal due to energy exchange between different modes. When $V(t)$ is modulated at twice the plasma frequency, the system undergoes parametric amplification, where a signal at frequency $f_{\mathrm{d}}/2+\delta f$ is amplified and an idler at frequency $f_{\mathrm{d}}/2-\delta f$ is created, with $\delta f$ being the tuning parameter, see Fig. 1(c). In Fig. 2, we demonstrate the parametric enhancement for the signal at $f_{\mathrm{s}}=30\,\mathrm{Hz}$ with respect to the system without pump. The system quickly reaches the steady state after 232-3 signal cycles and the enhancement persists over long times as required for robust amplification. The power spectrum exemplifies the amplification in frequency space.
@@
Figure 4: JPA characteristics. (a) Dimensionless gain $G_{\mathrm{s}}=S_{\omega}(A_{\mathrm{d}})/S_{\omega}(0)-1$ as a function of $A_{\mathrm{d}}$ is shown for signal amplitudes between $\tilde{I}_{1}=0.12$ and $0.42$. The results of the RSJ model are indicated by the dashed lines. (b) Signal-to-noise ratio (SNR) corresponding to the results shown in (a). (c, d) Simulation and RSJ model results at temperature $T/T_{0}=0.3$, using the same relative signal frequency with respect to the Josephson plasma frequency as in panels (a, b); see text.
Characterization of the JPA dynamics.— To characterize the dynamical regimes we map out the imbalance spectrum for signal frequencies between $f_{\mathrm{s}}=20$ and $100\,\mathrm{Hz}$, which exhibits JPA resonances and other excitation peaks  [^1] See Supplemental Material for the driven junction spectrum, the current-chemical potential characteristic, and the driven RSJ circuit model, which includes Refs. [54, 66].. Fig. 3 highlights the interplay between the signal and idler modes for the lowest JPA resonance. The mode weights are controlled by the pump amplitude. To describe the JPA dynamics we introduce a driven RSJ (resistively shunted junction) circuit model. The barrier velocity results in a modulation of the current across the junction given by $I_{1}\cos(\omega_{s}t)$, where $I_{1}=v_{1}I_{c}/v_{c}$ in which $v_{c}$ is the critical velocity associated with the critical current $I_{c}$ [65]. The barrier height modulation changes the tunneling energy, which can be incorporated as $I_{c}\bigl(1+A_{\mathrm{d}}\sin(\omega_{\mathrm{d}}t)\bigr)$. The Kirchhoff law of the driven RSJ circuit reads
$I_{1}\cos(\omega_{s}t)=I_{c}\bigl(1+A_{\mathrm{d}}\sin(\omega_{\mathrm{d}}t)\bigr)\sin\phi-G\Delta\mu,$ (4)
where $\phi=\phi_{L}-\phi_{R}$ is the phase difference across the junction and $G$ is the conductance [65]. The Josephson relation for the phase dynamics is $\hbar\dot{\phi}=-\Delta\mu$, with $\Delta\mu$ playing the role of the voltage across the junction. We numerically solve Eq. 4 for various signal and pump parameters and obtain the time evolution $\Delta\mu(t)$. The values of $I_{c}=135\,(10^{3}/\mathrm{s})$ and $\hbar G=59.8$ are determined from the simulated current-chemical potential relation [65]. With $I_{c}$ and the effective charging energy $E_{C}=4(\partial\mu/\partial N)$ we estimate the Josephson frequency $\omega_{\text{J}}=\sqrt{I_{c}E_{C}/\hbar}$ and obtain $f_{\text{J}}=38\,\mathrm{Hz}$ in agreement with the value of $f_{0}$. The result of the spectrum $|\Delta\mu(\omega)|^{2}$ is shown in Fig. 3d, which captures the dynamical response of the JPA. We note that the circuit model also describes higher JPA resonance and other excitation peaks well [65].
Figure 5: Dephasing dynamics. (a) Average vortex number $N_{v}$ as a function of $A_{\mathrm{d}}$ for the parameters used in Fig. 4(a). (b, c) Time evolution of the junction phase $\phi(t)$ for two different samples at $\tilde{I}_{1}=0.42$ with (continuous line) and without pump (dashed line). (d) Imbalance $z(t)$, averaged over many samples, corresponding to the parameters in (b, c).
To quantify the overall amplification, we define a dimensionless gain $G_{\mathrm{s}}=S_{\omega}(A_{\mathrm{d}})/S_{\omega}(0)-1$, where $S_{\omega}(A_{\mathrm{d}})$ represents the signal peak value and $S_{\omega}(0)$ corresponds to the result without pump. In Fig. 4(a) we demonstrate that $G_{\mathrm{s}}$ initially increases approximately quadratically with $A_{\mathrm{d}}$, and then reaches a maximum for a value of $A_{\mathrm{d}}$, that decreases with increasing $\tilde{I}_{1}=I_{1}/I_{c}$. For larger values of $A_{\mathrm{d}}$, the gain $G_{\mathrm{s}}$ decreases due to the dephasing dynamics as explained in the subsequent discussion. While the nonlinear increase is captured by the circuit model, the model fails to describe the decrease beyond the maximum, primarily visible for larger $\tilde{I}_{1}$. Fig. 4(b) presents the signal-to-noise ratio (SNR), defined as $\bar{S}_{\omega}/\Delta\bar{S}_{\omega}$, where we compute the mean signal value $\bar{S}_{\omega}=\left\langle S_{\omega}(A_{\mathrm{d}})-S_{\omega}(0)\right\rangle$ and the noise $\Delta\bar{S}_{\omega}$ from the variance determined by $\Delta\bar{S}_{\omega}^{2}=\left\langle(S_{\omega}(A_{\mathrm{d}})-S_{\omega}(0)-\bar{S}_{\omega})^{2}\right\rangle$. Here, $\langle..\rangle$ denotes a statistical average over $100$ samples of the ensemble. The SNR depends on both $A_{\mathrm{d}}$ and $\tilde{I}_{1}$, with a visible increase with increasing $A_{\mathrm{d}}$. The maximal regime of the SNR is in the same regime as the maximal gain $G_{\mathrm{s}}$, suggesting that this is the optimal regime of operation. In Fig. 4(c, d), the simulation results obtained at a higher temperature of $T/T_{0}=0.3$ show a noticeable reduction in overall gain, while still exhibiting a qualitatively similar trend to that observed at $T/T_{0}=0.06$. Thermal fluctuations suppress the Josephson plasma frequency $f_{0}$ and the critical current $I_{c}$, resulting in $f_{0}=35.2\,\mathrm{Hz}$ and $I_{c}=103\,(10^{3}/\mathrm{s})$ for $T/T_{0}=0.3$, see End Matter and [65] for details.
@@
Figure 6: Underlying nonlinear mechanism. (a) Time evolution of the density $\delta n(x,t)=n({\bf r},t)-n({\bf r},0)$, averaged over the $y$ direction and the ensemble, is shown for $\tilde{I}_{1}=0.21$ without the pump. (b) Dynamics in the presence of pump with $A_{\mathrm{d}}=0.8$. (c) Ensemble averaged phase difference $\phi$ at the junction for the cases in (a) and (b). (d) The results of zeroth and first order expansion of Eq. 4 without the resistive term; see text.
Fig. 5(a) shows the average total number of vortices in the system, determined by calculating the local phase winding around each lattice plaquette and averaging over both the time evolution and the initial ensemble, see End Matter for details. The increase in vortex number coincides with the slowdown and eventual reduction of the gain observed in Fig. 4(a), confirming that vortex excitations are the primary dissipation mechanism limiting the amplification. During time evolution, the junction phase undergoes phase-slippage events, where the local phase difference at the barrier exceeds $\pi$, see Figs. 5(b, c). These events result in the creation of vortex-antivortex pairs near the junction, providing a dissipation channel for the injected pump energy. When averaged over many realizations, these stochastic slippage events give rise to dephasing in the imbalance dynamics, as shown in Fig. 5(c). This dephasing reflects the cumulative impact of vortex formation, which gradually suppresses coherent amplification and ultimately sets the limit for the achievable gain.
To understand the underlying nonlinear mechanism of the JPA we examine the time evolution of the density and the phase at the junction in Fig. 6. Without pump, the barrier velocity modulation results in a smooth excitation of the density, which is created by the change of junction phase that is synchronized with the signal frequency. When the pump field is on, there are higher-order density excitations that result from the modulation of the phase by the pump field. These can be explained with a perturbative solution of the equation $\tilde{I}_{1}\cos(\omega_{\mathrm{s}}t)=\bigl(1+A_{\mathrm{d}}\sin(\omega_{\mathrm{d}}t)\bigr)\sin\phi$, assuming a small modulation parameter $A_{\mathrm{d}}$ and expanding $\phi$ as $\phi=\phi_{0}+A_{\mathrm{d}}\phi_{1}+\mathcal{O}(A_{\mathrm{d}}^{2})$. At zeroth order we obtain $\phi_{0}\approx\tilde{I}_{1}\cos(\omega_{\mathrm{s}}t)$, which explains the smooth oscillation of the phase without the pump. The first-order term gives $\phi_{1}\approx-\tilde{I}_{1}\cos(\omega_{\mathrm{s}}t)\sin(\omega_{\mathrm{d}}t)\bigr)$, highlighting the intrinsic nonlinear mixing between the pump and signal field. In Fig. 6(d), the perturbative solutions consistently capture the dynamics of our atomic JPA.
@@
Conclusions and outlook.— We have studied the dynamics of a driven atomic Josephson junction (JJ) using classical-field simulations. The JJ is formed by coupling two two-dimensional atom clouds with a tunnel barrier, where we periodically modulate both the position (input signal) and the height (pump) of the barrier. By tuning the pump frequency to twice the Josephson plasma frequency, we have demonstrated parametric amplification for the signal. Our system works in the $\mathrm{Hz}$ frequency range, that is beyond the operating range of the superconducting JPAs, thereby offering potential for low-frequency signal detection and precision sensing of slowly varying fields, such as gravitational or magnetic fields [Sheng2017, Javor2021, Somiya2023]. The current cold-atoms technology allows to accurately image the microscopic dynamics of the system [57, 58]. In this case, we have an unprecedented access to the density and phase microscopic dynamics at the junction responsible for the amplification and the intrinsic nonlinear mixing between the signal and pump fields. Our results can be experimentally probed using in-situ imaging and matter-wave interferometry, where the former enables measurement of the density imbalance and the latter allows detection of the underlying JPA mechanism. Our proposed protocol can be implemented in driven Josephson junctions experimentally realized in [57, 58].
@@
Our findings have broad implications for the understanding of the basic mechanism on which parametric amplification relies in quantum systems. For the specific case of quantum electronics devices, while the microscopic phase-density dynamics can be inferred through suitable Josephson current measurements, their direct experimental probing is out of the reach of current techniques, such as scanning probe microscopy.